Fill in the blanks to make the following statements true:
 (–5) + (– 8) = (– 8) + (............)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us assume the missing integer be x,
Then,
= (–5) + (–8) = (–8) + (x)
= - 5–8 = - 8 +x
= - 13 = - 8 + x
By sending–8 from RHS to LHS it becomes 8,
= - 13 + 8 = x
= x = - 5
Now substitute the x value in the blank place,
(–5) + (–8) = (–8) + (- 5)... [This equation is in the form of Commutative law ofAddition]

Thanks